Circular No. 6144

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA IN THE LARGE MAGELLANIC CLOUD 1995 M. Della Valle and N. Masetti, Department of Astronomy, University

of Padua; and S. Benetti, European Southern Observatory (ESO), report: “Preliminary inspection of a fully-reduced CCD spectrogram (range 380– 900 nm, resolution ∼ 1 nm), obtained on Mar. 3.03 UT with the ESO New Technology Telescope (+ EMMI), confirms Liller’s object (cf. IAUC 6143) to be a nova at or just past maximum light. The spectrum is dominated by strong emission lines of H, Fe II (multiplets 27, 37, 38, 42, 48, 49, 74), Ca II (H and K, and infrared triplet), and Na I (D lines and multiplet 5) flanked by narrow P-Cyg absorptions superimposed on a blue continuum. Additional features are observed at 553.7 nm [a blend of N II (multiplet 63) and Fe II (multiplet 55)] and at 747.6 and 778.0 nm [O I (multiplets 1 and 55, respectively)]. The expansion velocity measured from the minimum of the P-Cyg absorptions of the strongest lines is 900 ± 100 km/s. Hα and Hβ exhibit a blue P-Cyg absorption at ∼ 1500 km/s.”

P. M. Kilmartin and A. C. Gilmore provide a position for this nova, measured from an astrograph film taken at Mount John University Observatory on Mar. 3.4 UT: α = 5^{h}26^{m}50^{s}.33, δ = −70^{o}01′23′′.8 (equinox 2000.0; utilized 8 PPM stars with average residuals 0′′.3 in both α and δ).

SUPERNOVA 1995D IN NGC 2962 R. Tuffs, Max-Planck-Institut für Kernphysik, Heidelberg; A. Boselli,

Observatoire de Paris; G. Gavazzi, Osservatorio Milano; and H. Hippelein, Max-Planck-Institut für Astronomie, Heidelberg, report the following infrared magnitudes, obtained on Feb. 15.99 UT using the Calar Alto 2.2-m telescope (+ MAGIC camera): H = 14.1 ± 0.1, K = 14.0 ± 0.1.

Additional magnitudes: Feb. 19.07 UT, R = 13.2 ± 0.1 (M. Blake and T. Fuller, St. Mary’s University; comparison stars GSC 0242.0194, 0242.0810); Feb. 20.95, V = 13.40 ± 0.02, B−V = +0.22 ± 0.03 (H. Mikuž, Ljubljana; cf. IAUC 6137); Feb. 28.92, V = 13.64 ± 0.02, B −V = +0.41 ± 0.03 (Mikuž); Feb. 28.943, 13.44 (S. Meister, Bulach, Switzerland; unfiltered CCD; comparison star GSC 0242.0494); Mar. 1.915, 13.51 (Meister).

SUPERNOVA 1995E IN NGC 2441 Mikuž reports the following CCD photometry: Feb. 28.84 UT, V =

16.25 ± 0.03, B − V = +0.99 ± 0.05 (comparison star HR 3075).

1995 March 3 Daniel W. E. Green